# Improving Matrix Exponential for Generative AI Flows: A Taylor-Based Approach Beyond Paterson–Stockmeyer

### v1

@@ 1 Introduction
A significant breakthrough in this area was the development of advanced matrix polynomial evaluation techniques [9] that surpass the classical Paterson–Stockmeyer method [13]. These techniques, predicated on structured multiplications and linear combinations of matrix polynomials, enable higher-order Taylor approximations with reduced computational effort. For instance, the method presented in [22] introduced $m+$ approximations, where the first $m$ terms match the Taylor series exactly, while additional terms are optimized to enhance accuracy. These schemes, such as the $m=15+$ case introduced in [9, Ex. 5.1], achieve greater precision than standard Taylor polynomials of equivalent computational cost. Furthermore, recent optimizations for determining evaluation formula coefficients [18] have shown promising results for broader matrix function computations [18, 19].
The cost of these polynomial approximation methods is dominated by matrix multiplications, denoted by $M$. In practice, these operations scale as $\mathcal{O}(n^{3})$ for $n\times n$ matrices, whereas matrix additions and scalar multiplications are of order $\mathcal{O}(n^{2})$. In rational methods, matrix inverses are computed by solving linear systems, the cost of which can also be expressed in terms of matrix products as [12, App. C]:
@@
Table 1 compares the cost and achievable approximation orders of various state-of-the-art methods, including Padé approximants, Paterson–Stockmeyer evaluation, and mixed rational-polynomial schemes [11, 22, 14]. The results highlight the superior efficiency of modern polynomial evaluation strategies, which facilitate higher-order approximations using fewer matrix multiplications.
In this paper, we propose an optimized implementation of the matrix exponential based on the evaluation methods described in [9]. Our approach is specifically tailored for the high-throughput requirements of Generative AI flows, where the repeated evaluation of the matrix exponential often constitutes a primary computational bottleneck. Unlike traditional software routines optimized for fixed IEEE precision using pre-computed thresholds [21, 14, 22, 15], or flexible implementations that depend on external specialized libraries [16], we introduce a self-contained framework that dynamically selects the Taylor order and scaling factor for any user-defined tolerance $\varepsilon$. While previous dynamic methods have utilized classical Paterson–Stockmeyer schemes with more complex scaling strategies [17], the proposed algorithm leverages higher-order polynomial evaluations with a simplified operational logic. This provides a portable, library-independent solution that balances numerical stability with significant reductions in execution time for large-scale generative models.
@@
Approx. order from method [11, Tab. 3] 9 12 16 21 28
Padé evaluation cost $3.33M$ $4.33M$ $5.33M$ $6.33M$ $7.33M$
Approx. order Padé method [23, Tab. 2.2] 6 10 14 18 26
Table 1: Cost and available order of approximation with the polynomial evaluation methods from Paterson–Stockmeyer [13], Sastre–Ibáñez–Defez [22, Sec. 3], Bader–Blanes–Casas [14, Sec. 4], the mixed rational and polynomial approximation method from Sastre [11, Tab. 3] and the Padé approximation method from Higham [23, Tab. 2.2].
@@ 2.2 Generative flows
One approach [25] proposed replacing each weight matrix $W_{i}$ with its matrix exponential $e^{W_{i}}$. This guarantees invertibility, improves expressiveness over triangular matrices, and enables efficient computation of both the inverse and the log-determinant. Specifically, the inverse is given by $(e^{W})^{-1}=e^{-W}$, and the log-determinant simplifies to $\log\det(e^{W})=\operatorname{Tr}(W)$, reducing the computational cost from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
@@
For the sake of completeness the algorithm for the computation of the matrix exponential described in [25, Sec. 3.2] is reproduced here as Algorithm 1. It relies on the matrix exponential Taylor series expansion [2, Sec. 3], denoted by $T_{m}(W)$ and given by
@@
Note that the truncated Taylor series is particularly well-suited for scenarios with limited computational resources, restricted access to external libraries or large-scale problems involving well-behaved matrices [26, p. 52]. According to [25, Sec. 3.2], the weight matrices in neural networks, denoted by $W$, are often small in norm such that $|W|_{1}$ is also small. This property makes the incorporation of the matrix exponential into neural network architectures computationally feasible. Exploiting the identity $e^{W}=(e^{W/2^{s}})^{2^{s}}$, in Algorithm 1 the matrix is first scaled to ensure that $|W/2^{s}|<1/2$, after which the exponential is computed and subsequently squared $2^{s}$ times. This scaling and squaring strategy further reduces the computational cost while maintaining numerical stability; see [25, Alg. 1]. The computational cost of evaluating a Taylor approximation of degree $m$ using [25, Alg. 1], denoted by $T_{m}(A)$, is measured in terms of matrix multiplications. Let $C_{\text{orig}}$ represent the cost of evaluating the unscaled Taylor polynomial in Algorithm [25, Alg. 1], which is given by
$C_{\text{orig}}=m-1.$ (7)
When scaling is applied, the total cost becomes (s+m1)M(s+m-1)M, where $s$ is the scaling parameter. As reported in [25, Sec. 6.4], a tolerance of $\varepsilon=10^{-8}$ is used. Under this setting, the total number of matrix products, given by s+m1s+m-1, does not exceed 11, with an average value of 9.28 observed during training [25, Tab. 6].
Cost (7) makes Algorithm 1 unsuitable for high-dimensional data [25, Sec. 3.2]. To address this limitation, a second method is proposed that combines matrix exponentials with neural networks, using a low-rank parameterization of the weight matrix $W=A_{1}A_{2}$, where $A_{1}\in\mathbb{R}^{n\times t}$ and $A_{2}\in\mathbb{R}^{t\times n}$.
@@
where the truncation error bound of the matrix series of $V$, denoted here by $R^{\prime}_{m}(V)$, is given by (8) from [25, Sec. 3.2] reproduced here
@@
The approximation method is similar to Algorithm 1, with modifications: set the scale coefficient $s:=0$, change line 4 to $Y:=W/2$, and line 5 to $k:=3$ [25, Sec. 3.2].
While the cost of the algorithms proposed in [25, Sec. 3.2] represent a reasonable computational cost, further reductions are possible by optimizing the polynomial evaluation strategy.
@@ 3 Proposed method
Our implementation is specifically tailored for this generative flow setting. It builds upon efficient polynomial evaluation techniques, extending beyond the classical Paterson–Stockmeyer method as introduced in [9], to substantially reduce the number of matrix products required for a specified approximation order. Furthermore, the algorithm incorporates a dynamic procedure that optimally selects the Taylor series degree ($m$) and the scaling factor ($s$) to minimize the overall computational cost. This selection ensures that the error tolerance is strictly maintained, chosen to be no less than the unit roundoff ($u$) of the arithmetic used, while also strategically reducing the risk of overscaling [21, Sec. 1].
@@ 3.1 Evaluation formulas of the Taylor-based matrix exponential approximations
    The cost of evaluating $T_{8}(A)$ using formulas (13) and (3.1) is $3M$: one multiplication for computing $A^{2}$, one for evaluating (13), and one for evaluating (3.1). In contrast, computing the Taylor approximation of degree $m=8$ using the original algorithm from [25, Alg. 1], as given by (7), requires $7M$, which is more than twice as expensive.
@@
    This value differs from the exact Taylor coefficient $1/16!$ by a relative error of approximately $|b_{16}-1/16!|\cdot 16!\approx 0.454$. Consequently, $y_{22}(A)$ provides a more accurate approximation than $T_{15}(A)$ in exact arithmetic, although it remains less accurate than the full Taylor approximation $T_{16}(A)$ as demonstrated in [22]. Approximations of this type are referred to in [22] as $T_{15+}(A)$. Similarly, in this case, the cost of evaluating $y_{22}(A)$ using formulas (15)-(17) is $4M$.
@@ 3.2 Error analysis
can be significantly strict, i.e., $\|A^{k}\|\ll\|A\|^{k}$, as discussed in [21, Sec. 1]. This property is exploited in [5, Th. 1.1] to derive tighter bounds for the matrix power series, ensuring efficiency without the overhead of overscaling. For completeness of the exposition, the theorem is reproduced below:
@@
Let $A\in\mathbb{C}^{n\times n}$ be a square matrix and let $m\in\mathbb{N}$. The remainder of the matrix exponential Taylor approximation of degree $m$ given by (5)
@@
$\sum_{k=m+1}^{\infty}\frac{x^{k}}{k!}=\tilde{h}_{m+1}(x)<\frac{x^{m+1}}{(m+1)!}\cdot\frac{1}{1-\frac{x}{m+2}},\quad\text{for }x<m+2.$ (30)
@@
We verify the fulfillment of condition (28) by utilizing the error tolerance $\varepsilon=10^{-8}$ reported in [25, Sec. 6.4] and the selected Taylor degrees $m\in\{1,2,4,8,15\}$ (as defined in Section 3.1). For these parameters, it is easy to show that the condition is satisfied since the maximum value of the bound, $\alpha_{p}/2^{s}$, meets the requirement
@@
The error analysis for (8) is analogous to the error analysis described here, and the following theorem can be proved similarly:
@@
Finally, regarding the approximation in (8), we set the scaling parameter to $s=0$, consistent with the approach described in [25, Sec. 3.2]. For Taylor based approximations of approximation orders $m\in\{1,2,4,8,15+\}$, we employ the efficient evaluation formulas detailed in Section 3.1. For degrees exceeding these values, we utilize the classical Paterson–Stockmeyer method.
@@ 3.3 Proposed algorithms
Taking into account the evaluation formulas from Section 3.1 and the error analysis from Section 3.2, Algorithm 2 computes the matrix exponential using the scaling and squaring algorithm and the Taylor polynomial approximation. Polynomials are efficiently evaluated by means of Paterson-Stockmeyer method or evaluation formulas (10)-(17). Error tolerance $\varepsilon$ is variable, but it will be fixed to $10^{-8}$ in the experiments.
@@
5:Evaluate the matrix polynomial $X=P_{m}(W)$ by Paterson-Stockmeyer technique or by evaluation formulas (10)-(17)
@@
The algorithm is fundamentally suitable for any error tolerance $\varepsilon$ that is greater than or equal to the unit roundoff error. However, note that when employing IEEE double precision arithmetic, where $u=2^{-53}\approx 1.11\times 10^{-16}\ll 10^{-8}$, to achieve a higher level of accuracy that approaches the machine precision the algorithm parameters should be adjusted. Specifically, the accuracy is improved by selecting more terms in the Taylor approximation as shown in [22, Sec. 6].
Initially, Algorithm 2 determines the optimal parameters $m$ and $s$ by invoking Algorithm 3 or Algorithm 4. Next, the matrix $W$ is scaled following the scaling and squaring approximation to reduce its norm. Then, the matrix polynomial is evaluated either using the Paterson–Stockmeyer method or evaluation formulas (10)-(17). Finally, the exponential of the matrix $W$ is recovered carrying out the appropriate $s$ matrix products corresponding to the squaring stage.
In the implemented functions derived from Algorithms 2, 3, and 4, log2 has been used to avoid the appearance of large numbers when computing high powers of $\alpha_{p}$ and factorials in expressions like (34), especially if low precision arithmetic is used.
@@
Algorithm 3 determines the polynomial order $m$ and the scaling parameter $s$ required to compute the matrix exponential of $W$ using a Taylor approximation of order $m\leq 16$ and a desired tolerance $\varepsilon$. It is assumed that the polynomials are evaluated using the Paterson–Stockmeyer method.
@@
17:  $p=2\cdot i-1$
18:  if $m=1$ then
19:   $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
20:   $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
21:  else
22:   $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
23:   $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}^{2}\|_{1}$
24:  end if
25:  if $E_{1}+E_{2}\leq\varepsilon$ then
26:   $finish=1$
27:  else
28:   $i=i+1$
29:  end if
30:end while
31:if $finish=0$ then
32:  for $i=1$ to 22 do
33:   $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
34:   if $s1>s$ then
35:     $s=s1$
36:   end if
37:  end for
38:end if
39:if $s>20$ then
40:  $s=20$
@@
Next, from lines 9 to 28, an attempt is made to determine the most appropriate degree $m$ of the polynomial to use, from the point of view of accuracy and computational efficiency. Since $\frac{W^{m+1}}{(m+1)!}$ and $\frac{W^{m+2}}{(m+2)!}$ are the first two terms of the remainder for the Taylor-based approximation of order $m$ to the exponential of $W$, our objective is to obtain the value of $m$ that guarantees that:
@@
Notwithstanding, since $W^{m+1}$ and $W^{m+2}$ are not explicitly calculated, the 1-norm of these matrix powers is bounded using the 1-norm of the highest power of $W$ computed thus far. For example, when $m=1$ (lines 16–18), $W$ is known, but $\|W^{2}\|_{1}$ and $\|W^{3}\|_{1}$ are required; these are bounded by $\|W\|_{1}^{2}$ and $\|W\|_{1}^{3}$, respectively. Similarly, if $m=12$ (lines 19–21), $W^{4}$ has been previously calculated, and the required norms $\|W^{13}\|_{1}$ and $\|W^{14}\|_{1}$ are bounded by $\|W^{4}\|_{1}^{3}\cdot\|W\|_{1}$ and $\|W^{4}\|_{1}^{3}\cdot\|W^{2}\|_{1}$. Thus, if $m$ equals 12, $j$ and $k$ take the values 4 and 3, accordingly.
@@
in accordance with lines 30 to 36 in the algorithm. Finally, to avoid overscaling, the maximum value of $s$ is set to 20 in lines 37 and 38.
Similarly, Algorithm 3 fulfills the same role as Algorithm 2, but incorporates formulas (10)-(17) for evaluating the matrix polynomials. While the structure remains consistent with the previous algorithm, the polynomial degrees differ. Furthermore, the maximum explicitly computed power of $W$ is restricted to $W^{2}$. A key distinction lies in the penultimate value of vector $C$, which corresponds to the Taylor-based approximation $T_{15+}$ defined in (17). As previously noted, this approximation is equivalent to (18), where $b_{16}$ is given in IEEE double-precision arithmetic by (20). Consequently, the penultimate element of vector $C$ is derived from the corresponding remainder term (19). For $m=15$, the indices $j$ and $k$ are set to 2 and 8, respectively. Thus, using $W^{2}$, the bounds for $\|W^{16}\|_{1}$ and $\|W^{17}\|_{1}$ (lines 20, 21, and 26) are bounded by $\|W^{2}\|_{1}^{8}$ and $\|W^{2}\|_{1}^{8}\cdot\|W\|_{1}$. The calculation of the scaling parameter $s$ follows the procedure described in Algorithm 2. However, in this case determining $s$ based on the maximum value from the first two terms of the remainder (19) ensures a more conservative selection.
Algorithm 4 Algorithm for Obtaining an Polynomial Order $m$ and a Scaling Parameter $s$. The polynomial will be evaluated via evaluation formulas (10)-(17)

### v2

@@ 1 Introduction
A significant breakthrough in this area was the development of advanced matrix polynomial evaluation techniques [8] that surpass the classical Paterson–Stockmeyer method [12]. These techniques, predicated on structured multiplications and linear combinations of matrix polynomials, enable higher-order Taylor approximations with reduced computational effort. For instance, the method presented in [20] introduced $m+$ approximations, where the first $m$ terms match the Taylor series exactly, while additional terms are optimized to enhance accuracy. These schemes, such as the $m=15+$ case introduced in [8, Ex. 5.1], achieve greater precision than standard Taylor polynomials of equivalent computational cost. Furthermore, recent optimizations for determining the coefficients of evaluation formulas [17] have shown promising results for broader matrix function computations [17, 18].
The cost of these polynomial approximation methods is dominated by matrix multiplications, denoted by $M$. In practice, these operations scale as $\mathcal{O}(n^{3})$ for $n\times n$ matrices, whereas matrix additions and scalar multiplications are of order $\mathcal{O}(n^{2})$. In rational methods, matrix inverses are computed by solving linear systems, the cost of which can also be expressed in terms of matrix products as [11, App. C]:
@@
Table 1 compares the cost and achievable approximation orders of various state-of-the-art methods, including Padé approximants [21], Paterson–Stockmeyer evaluation [12], advanced matrix polynomial evaluation [8, 20, 13] and mixed rational-polynomial schemes [10]. The results highlight the superior efficiency of modern polynomial evaluation strategies, which facilitate higher-order approximations using fewer matrix multiplications.
In this paper, we propose an optimized implementation of the matrix exponential based on the matrix polynomial evaluation methods described in [8]. Our approach is specifically tailored for the high-throughput requirements of Generative AI flows, where the repeated evaluation of the matrix exponential often constitutes a primary computational bottleneck. Unlike traditional software routines optimized for fixed IEEE precision using pre-computed thresholds [19, 13, 20, 14], or flexible implementations that depend on external specialized libraries [15], we introduce a self-contained framework that dynamically selects the Taylor order and scaling factor for any user-defined tolerance $\varepsilon$. While previous dynamic methods have utilized classical Paterson–Stockmeyer schemes with more complex scaling strategies [16], the proposed algorithm leverages higher-order polynomial evaluations with a simplified operational logic. This provides a portable, library-independent solution that balances numerical stability with significant reductions in execution time for large-scale generative models.
@@
Approx. order from method [10, Tab. 3] 9 12 16 21 28
Table 1: Cost and available order of approximation with the polynomial evaluation methods from Paterson–Stockmeyer [12], Bader–Blanes–Casas [13, Sec. 4], Sastre–Ibáñez–Defez [20, Sec. 3], the Padé approximation method from Higham [21, Tab. 2.2] and the mixed rational and polynomial approximation method from Sastre [10, Tab. 3].
@@ 2.2 Generative flows
One approach [23] proposed replacing each weight matrix $W_{i}$ with its matrix exponential $e^{W_{i}}$. This guarantees invertibility, improves expressiveness over triangular matrices, and enables efficient computation of both the inverse and the log-determinant. Specifically, the inverse is given by $(e^{W})^{-1}=e^{-W}$, and the log-determinant simplifies to $\log|(e^{W})|=\operatorname{Tr}(W)$, reducing the computational cost from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
@@
For the sake of completeness the algorithm for the computation of the matrix exponential described in [23, Sec. 3.2] is reproduced here as Algorithm 1. It relies on the matrix exponential Taylor series expansion [2, Sec. 3], denoted by $T_{m}(W)$ and given by
@@
Note that the truncated Taylor series is particularly well-suited for scenarios with limited computational resources, restricted access to external libraries or large-scale problems involving well-behaved matrices [24, p. 52]. According to [23, Sec. 3.2], the weight matrices in neural networks, denoted by $W$, are often small in norm such that $\|W\|_{1}$ is also small. This property makes the incorporation of the matrix exponential into neural network architectures computationally feasible. Exploiting the identity $e^{W}=(e^{W/2^{s}})^{2^{s}}$, in Algorithm 1 the matrix is first scaled to ensure that $|W/2^{s}|<1/2$, after which the exponential is computed and subsequently squared $2^{s}$ times. This scaling and squaring strategy further reduces the computational cost while maintaining numerical stability; see [23, Alg. 1]. The computational cost of evaluating the Taylor approximation of degree $m$ from (4) using [23, Alg. 1] is measured in terms of matrix multiplications $(M)$. Let $C_{\text{orig}}$ represent the cost of evaluating the unscaled Taylor polynomial in Algorithm [23, Alg. 1], which is given by
$C_{\text{orig}}=(m-1)M.$ (7)
When scaling is applied, the total cost becomes (s+m1)M(s+m-1)M, where $s$ is the scaling parameter. As reported in [23, Sec. 6.4], a tolerance of $\varepsilon=10^{-8}$ is used. Under this setting, the total number of matrix products, given by s+m1s+m-1, does not exceed 11, with an average value of 9.28 observed during training [23, Tab. 6].
Cost (7) makes Algorithm 1 unsuitable for high-dimensional data [23, Sec. 3.2]. To address this limitation, a second method is proposed that combines matrix exponentials with neural networks, using a low-rank parameterization of the weight matrix $W=A_{1}A_{2}$, where $A_{1}\in\mathbb{R}^{n\times t}$ and $A_{2}\in\mathbb{R}^{t\times n}$.
@@
where the truncation error bound of the matrix series of $V$, denoted here by $R^{\prime}_{m}(V)$, is given by (8) from [23, Sec. 3.2] reproduced here
@@
The approximation method is similar to Algorithm 1, with modifications: set the scale coefficient $s:=0$, change line 4 to $Y:=W/2$, and line 5 to $k:=3$ [23, Sec. 3.2].
While the cost of the algorithms proposed in [23, Sec. 3.2] represent a reasonable computational cost, further reductions are possible by optimizing the polynomial evaluation strategy.
@@ 3 Proposed method
Our implementation is specifically tailored for this generative flow setting. It builds upon efficient polynomial evaluation techniques, extending beyond the classical Paterson–Stockmeyer method as introduced in [8], to substantially reduce the number of matrix products required for a specified approximation order. Furthermore, the algorithm incorporates a dynamic procedure that optimally selects the Taylor series degree ($m$) and the scaling factor ($s$) to minimize the overall computational cost. This selection ensures that the error tolerance is maintained, chosen to be no less than the unit roundoff ($u$) of the arithmetic used, while also strategically reducing the risk of overscaling [19, Sec. 1].
@@ 3.1 Evaluation formulas of the Taylor-based matrix exponential approximations
    The cost of evaluating $T_{8}(A)$ using formulas (13) and (3.1) is $3M$: one multiplication for computing $A^{2}$, one for evaluating (13), and one for evaluating (3.1). In contrast, computing the Taylor approximation of degree $m=8$ using the original algorithm from [23, Alg. 1], as given by (7), requires $7M$, which is more than twice as expensive.
@@
    This value differs from the exact Taylor coefficient $1/16!$ by a relative error of approximately $|b_{16}-1/16!|\cdot 16!\approx 0.454$. Consequently, $y_{22}(A)$ provides a more accurate approximation than $T_{15}(A)$ in exact arithmetic, although it remains less accurate than the full Taylor approximation $T_{16}(A)$ as demonstrated in [20]. Approximations of this type are referred to in [20] as $T_{15+}(A)$. Similarly, the evaluation of $y_{22}(A)$ via formulas (15)–(17) requires only $4M$ matrix products. This cost contrast is even more pronounced when using the low-rank parameterization from (8). Since no scaling is applied in that case, the evaluation of the corresponding matrix polynomial of degree 15 requires $14M$ products. Consequently, the baseline approach is 3.5 times more expensive than our proposed method in this configuration.
@@ 3.2 Error analysis
can be significantly strict, i.e., $\|A^{k}\|\ll\|A\|^{k}$, as discussed in [19, Sec. 1]. This property is exploited in [5, Th. 1.1] to derive tighter bounds for the matrix power series, ensuring efficiency without the overhead of overscaling. For completeness of the exposition, the theorem is reproduced below:
@@
Let $A\in\mathbb{C}^{n\times n}$ be a square matrix and let $m\in\mathbb{N}$. The remainder of the matrix exponential Taylor approximation of degree $m$ is given by (5)
@@
$\sum_{k=m+1}^{\infty}\frac{x^{k}}{k!}=\tilde{h}_{m+1}(x)<\frac{x^{m+1}}{(m+1)!}\cdot\frac{1}{1-\frac{x}{m+2}},\quad\text{for }0<x<m+2.$ (30)
@@
We verify the fulfillment of condition (28) by utilizing the error tolerance $\varepsilon=10^{-8}$ reported in [23, Sec. 6.4] and the selected Taylor degrees $m\in\{1,2,4,8,15\}$ (as defined in Section 3.1). For these parameters, it is easy to show that the condition is satisfied since the maximum value of the bound, $\alpha_{p}/2^{s}$, meets the requirement
@@
The error analysis for the series in (8) is analogous to the error analysis described here, and the following theorem can be proved similarly:
@@
Remark 1.
The proposed approach offers computational advantages for nonnormal matrices. Consider the well-known test matrix (1.2) from [19], defined as
$W=\begin{bmatrix}1&b\\ 0&-1\end{bmatrix}$,
taking for instance $b=10$. For this matrix, the baseline Algorithm 1 selects a scaling parameter $s=5$, whereas it is easy to show that our approach, based on Theorem 2, yields $s=0$. This reduction eliminates 5 matrix-matrix multiplications, demonstrating that our method effectively prevents the over-scaling typically induced by large off-diagonal entries. Furthermore, unlike the approach in [19], this is achieved without estimating norms of matrix powers. Such estimations incur an $\mathcal{O}(n^{2})$ overhead that remains non-negligible for moderate-sized matrices, as noted in [13] for $n\leq 150$ and [20, p. 213] for $n=128$.
Finally, regarding the approximation in (8), we set the scaling parameter to $s=0$, consistent with the approach described in [23, Sec. 3.2]. For Taylor based approximations of approximation orders $m\in\{1,2,4,8,15+\}$, we employ the efficient evaluation formulas detailed in Section 3.1. For degrees exceeding these values, we utilize the classical Paterson–Stockmeyer method.
@@ 3.3 Proposed algorithms
Taking into account the evaluation formulas from Section 3.1 and the error analysis from Section 3.2, Algorithm 2 computes the matrix exponential using the scaling and squaring algorithm and the Taylor polynomial approximation. Polynomials are efficiently evaluated by means of Paterson–Stockmeyer method or evaluation formulas (10)-(17). Error tolerance $\varepsilon$ is variable, but it will be fixed to $10^{-8}$ in the experiments [23, Sec. 6.4].
@@
5:Evaluate the matrix polynomial $X=P_{m}(W)$ by Paterson–Stockmeyer technique or by evaluation formulas (10)-(17)
@@
The algorithm is fundamentally suitable for any error tolerance $\varepsilon$ that is greater than or equal to the unit roundoff error. However, note that when employing IEEE double precision arithmetic, where $u=2^{-53}\approx 1.11\times 10^{-16}\ll 10^{-8}$, to achieve a higher level of accuracy that approaches the machine precision the algorithm parameters should be adjusted. Specifically, the accuracy is improved by selecting more terms in the Taylor approximation as shown in [20, Sec. 6].
Initially, Algorithm 2 determines the optimal parameters $m$ and $s$ by invoking Algorithm 3.3 or Algorithm 3.3. Next, the matrix $W$ is scaled following the scaling and squaring approximation to reduce its norm. Then, the matrix polynomial is evaluated either using the Paterson–Stockmeyer method or evaluation formulas (10)-(17). Finally, the exponential of the matrix $W$ is recovered carrying out the appropriate $s$ matrix products corresponding to the squaring stage.
In the implemented functions derived from Algorithms 2, 3.3, and 3.3, log2 has been used to avoid the appearance of large numbers when computing high powers of $\alpha_{p}$ and factorials in expressions like (34), especially if low precision arithmetic is used.
@@
Algorithm 3.3 determines the polynomial order $m$ and the scaling parameter $s$ required to compute the matrix exponential of $W$ using a Taylor approximation of order $m\leq 16$ and a desired tolerance $\varepsilon$. It is assumed that the polynomials are evaluated using the Paterson–Stockmeyer method.
@@
17:  if $j>k$ then
18:    $W^{j}=W^{j-1}\cdot W$
19:  end if
20:  $p=2\cdot i-1$
21:  if $m=1$ then
22:    $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
23:    $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
24:  else
25:    $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
26:    $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}^{2}\|_{1}$
27:  end if
28:  if $E_{1}+E_{2}\leq\varepsilon$ then
29:    $finish=1$
30:  else
31:    $i=i+1$
32:  end if
33:end while
34:if $finish=0$ then
35:  for $i=1$ to 22 do
36:    $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
37:    if $s1>s$ then
38:     $s=s1$
39:    end if
40:  end for
@@
Next, from lines 9 to 31, an attempt is made to determine the most appropriate degree $m$ of the polynomial to use, from the point of view of accuracy and computational efficiency. Since $\frac{W^{m+1}}{(m+1)!}$ and $\frac{W^{m+2}}{(m+2)!}$ are the first two terms of the remainder for the Taylor-based approximation of order $m$ to the exponential of $W$, our objective is to obtain the value of $m$ that guarantees that:
@@
Notwithstanding, since $W^{m+1}$ and $W^{m+2}$ are not explicitly calculated, the 1-norm of these matrix powers is bounded using the 1-norm of the highest power of $W$ computed thus far. For example, when $m=1$ (lines 19–21), just $W$ is known, but $\|W^{2}\|_{1}$ and $\|W^{3}\|_{1}$ are required; these are bounded by $\|W\|_{1}^{2}$ and $\|W\|_{1}^{3}$, respectively. Similarly, if, for example, $m=12$ (lines 22–24), $W^{4}$ will have been previously calculated, and the required norms $\|W^{13}\|_{1}$ and $\|W^{14}\|_{1}$ are bounded by $\|W^{4}\|_{1}^{3}\cdot\|W\|_{1}$ and $\|W^{4}\|_{1}^{3}\cdot\|W^{2}\|_{1}$. Thus, if $m$ equals 12, $j$ and $k$ take the values 4 and 3, accordingly. In detail, Table 4 contains the values of $j$ and $k$ used for any value of $m$ taken into account.
@@
in accordance with lines 32 to 39 in the algorithm. Finally, to avoid overscaling, the maximum value of $s$ is set to 20 in lines 40 and 41.
$M$ Computed $W$ Power Required Norms Estimated Norms
1 $W$ $\|W^{2}\|_{1}$ $\|W^{3}\|_{1}$ $\|W\|^{2}_{1}$ $\|W\|^{3}_{1}$
2 $W^{2}$ $\|W^{3}\|_{1}$ $\|W^{4}\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W^{2}\|_{1}$
4 $W^{2}$ $\|W^{5}\|_{1}$ $\|W^{6}\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W^{2}\|_{1}$
6 $W^{3}$ $\|W^{7}\|_{1}$ $\|W^{8}\|_{1}$ $\|W^{3}\|^{2}_{1}\cdot\|W\|_{1}$ $\|W^{3}\|^{2}_{1}\cdot\|W^{2}\|_{1}$
9 $W^{3}$ $\|W^{10}\|_{1}$ $\|W^{11}\|_{1}$ $\|W^{3}\|^{3}_{1}\cdot\|W\|_{1}$ $\|W^{3}\|^{3}_{1}\cdot\|W^{2}\|_{1}$
12 $W^{4}$ $\|W^{13}\|_{1}$ $\|W^{14}\|_{1}$ $\|W^{4}\|^{3}_{1}\cdot\|W\|_{1}$ $\|W^{4}\|^{3}_{1}\cdot\|W^{2}\|_{1}$
16 $W^{4}$ $\|W^{17}\|_{1}$ $\|W^{18}\|_{1}$ $\|W^{4}\|^{4}_{1}\cdot\|W\|_{1}$ $\|W^{4}\|^{4}_{1}\cdot\|W^{2}\|_{1}$
$m$ $W^{j}$ $\|W^{m+1}\|_{1}$ $\|W^{m+2}\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W^{2}\|_{1}$
Table 4: Computed powers of $W$, the corresponding norms required by (42), and their resulting approximations for each polynomial degree $M$ considered in Algorithm 3.3.
Similarly, Algorithm 3.3 fulfills the same role as Algorithm 3.3, but incorporates formulas (10)-(17) for evaluating the matrix polynomials. While the structure remains consistent with the previous algorithm, the polynomial degrees differ. Furthermore, the maximum explicitly computed power of $W$ is restricted to $W^{2}$. A key distinction lies in the penultimate value of vector $C$, which corresponds to the Taylor-based approximation $T_{15+}$ defined in (17). As previously noted, this approximation is equivalent to (18), where $b_{16}$ is given in IEEE double-precision arithmetic by (20). Consequently, the penultimate element of vector $C$ is derived from the corresponding remainder term (19). Table 5 incorporates the values of $j$ and $k$ applied for any value of $m$. For $m=15$, the indices $j$ and $k$ are set to 2 and 8, respectively. Thus, using $W^{2}$, the bounds for $\|W^{16}\|_{1}$ and $\|W^{17}\|_{1}$ (lines 22, 23, and 28) are bounded by $\|W^{2}\|_{1}^{8}$ and $\|W^{2}\|_{1}^{8}\cdot\|W\|_{1}$. The calculation of the scaling parameter $s$ follows the procedure described in Algorithm 3.3. However, determining $s$ based on the maximum value from the first two terms of the remainder (19) ensures in this case a more conservative selection.
Algorithm 4 Algorithm for Obtaining a Polynomial Order $m$ and a Scaling Parameter $s$. The polynomial will be evaluated via formulas (10)–(17)